OpenStudy (anonymous):

How do I solve problems like this? Solve 8 = 2^x + 4. please help me fan / medal

OpenStudy (anonymous):

Www.wolframalpha can tell you

hartnn (hartnn):

is it \(\Large 2^{x+4}\) ?

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

no its like hartan said

hartnn (hartnn):

then write \(\Large 8 = 2^3\) and compare the exponents!

OpenStudy (anonymous):

?

hartnn (hartnn):

\(\Large 2^3 = 2^{x+4} \implies 3 = x+4\) can you find x from here ?

OpenStudy (anonymous):

X=2 right?

OpenStudy (anonymous):

8 = 2-1+4

hartnn (hartnn):

no... 3 = x+4 subtract 4 from both sides what u get ?

OpenStudy (anonymous):

hola